The number of boys and girls in a class are in the ratio 7:5. The number of boys is 8 more than the number of girls. What is the total class strength?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the number of boys be 7x and girls be 5x.
According to the question, 7x=5x+8
⇒7x–5x=8
⇒2x=8
⇒x=8/2
⇒x=4
Therefore, Number of boys = 7×4 = 28 And,
Number of girls = 5×4 = 20
Total number of students = 20+28 = 48

Thanks